# Enhancing Remote Sensing Vision-Language Models for Zero-Shot Scene Classification thanks:

### v1

@@ Abstract
remote sensing, scene classification, vision-language models, zero-shot, transductive inference.
@@ I Introduction
Linking images with text descriptions has been demonstrated to be an effective approach for learning granular visual representations [7, 8]. While this idea seemed powerful, pioneering works in the field of RS [9, 10] were limited by computational budgets and the quantity of available RS data, both of which have been significant bottlenecks for generalization and robustness capabilities [11]. More recently, Vision-Language Models (VLMs) like CLIP [12] have overcome these limitations by leveraging a new pretraining paradigm that uses large-scale image-text pair datasets for unsupervised contrastive learning. These models have demonstrated high capability for numerous downstream tasks, including efficient zero-shot image classification by prompting arbitrary candidate class descriptions, e.g., "a satellite photo of a [class].", sometimes even surpassing supervised competitors [12]. Inspired by these promising results, the RS community has worked on developing large image-text RS datasets  [13, 14, 15, 16] leading to rapid progress in zero-shot scene classification benchmarks [17].
@@
In remote sensing scene classification, both the large size of the images and the need for granular information pose challenges. To make high-resolution inference tractable, it is common practice to divide the images into smaller patches and generate predictions for each patch individually; this is known as inductive inference. Another paradigm known as transductive inference [18, 19], has shown that jointly considering multiple instances at prediction time can improve the prediction accuracy by accounting for the statistical distribution of instances in the embedding space [20, 21].
Despite its large potential, transductive inference has been largely overlooked in RS within the context of VLMs. We aim to address this gap by introducing an efficient transductive method that operates exclusively within the embedding space, i.e., in a black box setup after feature extraction.
In a zero-shot classification setting, class-specific textual prompts are mapped to a shared embedding space generating individual pseudo-label for each image patch. In a traditional inductive inference process, predictions are generated by utilizing initial pseudo-labels to identify the most confident class, with each patch predicted individually. In contrast, our work envisions transductive inference in a zero-shot classification setting. This approach leverages the data structure within the feature space to account for instance relations, enabling collective prediction of all points simultaneously. Our proposed objective function can be viewed as a regularized maximum-likelihood estimation, constrained by a Kullback-Leibler divergence penalty that integrates the aforementioned initial pseudo-labels and a Laplacian term that constraints similar patches to have similar predictions.
@@ II-A Vision-Language Models for Remote Sensing
Due to foundation models like CLIP being trained on natural images, there is an active research effort to build domain-specific versions of these models. This is prevalent in the medical imaging community where VLMs have shown promising results [22, 23] in improving image-text retrieval and few-shot classification. The RS community has followed suit, working on creating extensive image-text datasets by scraping and filtering public satellite and UAV imagery sources [13, 14, 15, 16]. This has led to the development of several fine-tuned VLMs on various downstream tasks [24, 25, 26, 27, 28], with many of them showing strong performances in zero-shot scene classification  [11, 13, 15].
@@ II-B Transductive inference in Vision-Language Models
In the few-shot literature, transduction leverages both the few labeled samples and unlabeled test data outperforming inductive methods [29, 30, 31, 32]. However, when applied to VLMs, these transductive methods face significant performance drops [20, 21] since they are based solely on the vision features. This motivated very recent transductive methods in computer vision to explicitly leverage the textual modality alongside image embeddings – a capability not present before the emergence of VLMs [20, 21]. Building on these advances and the transductive-inference zero-shot objective described in [21], our work enhances the predictive accuracy of pretrained RS VLMs without the need of any supervision.
@@ III Method
The transductive approach employed by RS-TransCLIP is based on the hypothesis that the data structure within the feature space can be modeled as a mixture of Gaussian distributions. As a result, the RS-TransCLIP objective function integrates this hypothesis alongside affinity relationships among patches and initial text-based pseudo-labels to minimize prediction deviation.
@@ III-A Variable Definition
with $K$ the number of classes, $\mathcal{Q}$ the set of sample indices and $\Delta_{K}$ the $K$-dimensional probability simplex i.e. prediction space.
@@ III-B RS-TransCLIP objective function
GMM clustering. — The goal of this term is maximizing the similarity between the assignment variables ${\mathbf{z}}_{i}$ and the likelihood ${\mathbf{p}}_{i}$. In our case, we model the likelihood of target data as a balanced mixture of $K$ multivariate Gaussian distributions. Each distribution represents a class $k$ with an associated mean vector $\boldsymbol{\mu}_{k}$ and a covariance matrix $\boldsymbol{\Sigma}$. Defining ${\mathbf{f}}_{i}\in\mathbb{R}^{d}$ as the image embedding of sample $i$, we set $p_{i,k}$ the probability that sample $i$ is generated by the Gaussian distribution of class $k$:
@@
Laplacian regularization. — The aim of this term is to favor pairs of samples with high affinity to have similar assignment variables. In our case, we define non-negative affinities $w_{ij}$ using the cosine similarities between image embeddings of each sample (see line 1 in Algorithm 1). Note that affinities can be tailored for each specific use-case, provided the affinity matrix is positive semi-definite. This ensures the concavity of the term, which in turn guarantees the convergence of the decoupled updates (refer to [21] for details).
KL regularization. — The purpose of this term is to prevent the assignment variables to deviate significantly from the initial pseudo-labels. In our case, we obtain the pseudo-labels $(\hat{\mathbf{y}}_{i})_{1\leq i\leq\mathcal{Q}}$ by applying the softmax function to the vector whose components are obtained by computing the dot product between the image embeddings $\mathbf{f}_{i}$ and all text embeddings $\mathbf{t}_{k}\in\mathbb{R}^{d}$, scaled by the temperature factor $\tau$ used during VLM pretraining (see line 1 in Algorithm 1). This allows us to integrate the text-knowledge into the optimization process.
@@ IV-A Experimental setup
We test RS-TransCLIP on four VLMs: CLIP [12], RemoteCLIP [11], SkyCLIP [15], and GeoRSCLIP [13] — all with various model architectures to generate their respective image embeddings. Using RS text-prompt templates from [13], 106 individual text embeddings were averaged out to get a single textual embedding per class. The zero-shot scene classification performance is evaluated on 10 RS benchmark datasets: AID, EuroSAT, MLRSNet, OPTIMAL31, PatternNet, RESISC45, RSC11, RSICB128, RSICB256, and WHURS19 [33, 34, 35, 36, 37, 38, 39, 40, 41]. Note that none of the chosen VLMs were fine-tuned on any of the listed datasets. Table 1 presents the zero-shot top-1 accuracy, without and with the addition of RS-TransCLIP.
@@ IV-B Zero-shot classification — without RS-TransCLIP
Model
AID
EuroSAT
MLRSNet
OPTIMAL31
PatternNet
RESISC45
RSC11
RSICB128
RSICB256
WHURS19
Average
@@ V Conclusion
In this work, we proposed RS-TransCLIP, a transductive algorithm that enhances RS VLMs with minimal extra computational cost. By leveraging initial pseudo-labels and patch affinities, our method improves zero-shot capabilities through transductive inference, demonstrating significant accuracy improvements over inductive zero-shot classification and showing its wide applicability beyond natural images [21].
Future works will study RS-TransCLIP’s performance concerning text-prompt variability, given VLMs’ high sensitivity to input text prompts. Moreover, adapting RS-TransCLIP to a few-shot setting to incorporate labeled data will be explored in human-in-the-loop scenarios.

### v2

# Karim El Khoury*1      Maxime Zanella*1,2      Benoît Gérin*1      Tiffanie Godelaine*1 Benoît Macq1   Saïd Mahmoudi2   Christophe De Vleeschouwer1   Ismail Ben Ayed3 1UCLouvain, Belgium        2UMons, Belgium        3ÉTS Montreal, Canada

@@ Abstract
remote sensing, scene classification, vision-language models, zero-shot, transductive inference
@@ I Introduction
Linking images with text descriptions has been an effective approach for learning granular visual representations [7, 8]. While this idea seemed powerful, pioneering works in the field of RS [9, 10] were limited by computational budgets and the quantity of available RS data, both of which have been significant bottlenecks for generalization and robustness capabilities [11]. More recently, Vision-Language Models (VLMs) like CLIP [12] have overcome these limitations by leveraging a new pretraining paradigm that uses large-scale image-text pair datasets for unsupervised contrastive learning. These models have demonstrated high capability for numerous downstream tasks, including efficient zero-shot image classification by prompting arbitrary candidate class descriptions, e.g., "a satellite photo of a [class].", sometimes even surpassing supervised competitors [12]. Inspired by these promising results, the RS community has worked on developing large image-text RS datasets  [13, 14, 15, 16] leading to rapid progress in zero-shot scene classification benchmarks [17].
@@
In remote sensing scene classification, both the large size of the images and the need for granular information pose challenges. To make high-resolution inference tractable, it is common practice to divide the images into smaller patches and generate predictions for each patch individually; this is known as inductive inference. Another paradigm known as transductive inference [18, 19], has shown that jointly considering multiple instances at prediction time can improve the prediction accuracy by accounting for the statistical distribution of instances in the embedding space [20, 21]. Despite its large potential, transductive inference has been largely overlooked in RS within the context of VLMs. We aim to address this gap by introducing an efficient transductive method that operates exclusively within the embedding space, i.e., in a black box setup after feature extraction.
In a zero-shot classification setting, class-specific textual prompts are mapped to a shared embedding space generating individual pseudo-label for each image patch. In a traditional inductive inference process, predictions are generated by utilizing initial pseudo-labels to identify the most confident class, with each patch predicted individually. In contrast, our work envisions transductive inference in a zero-shot classification setting. As shown in Fig. 2, this approach leverages the data structure within the feature space to account for instance relations, enabling collective prediction of all points simultaneously. Our proposed objective function can be viewed as a regularized maximum-likelihood estimation, constrained by a Kullback-Leibler divergence penalty that integrates the aforementioned initial pseudo-labels and a Laplacian term that constraints similar patches to have similar predictions.
@@ II-A Vision-Language Models for Remote Sensing
Due to foundation models being trained on natural images, there is an active research effort to build domain-specific versions of these models. This is prevalent in the medical imaging where VLMs have shown promising results [22, 23] in improving image-text retrieval and few-shot classification. The RS community has followed suit, working on creating extensive image-text datasets by scraping and filtering public satellite and UAV imagery sources [13, 14, 15, 16]. This has led to the development of several fine-tuned VLMs on various downstream tasks [24, 25, 26, 27, 28], with many of them showing strong performances in zero-shot scene classification  [11, 13, 15].
@@ II-B Transductive inference in Vision-Language Models
In the few-shot literature, transduction leverages both the few labeled samples and unlabeled test data outperforming inductive methods [29, 30, 31, 32]. However, when applied to VLMs, these transductive methods face significant performance drops [20, 21] since they are based solely on the vision features. This motivated very recent transductive methods in computer vision to explicitly leverage the textual modality alongside image embeddings – a capability not present before the emergence of VLMs [20, 21, 33]. Building on these advances and the transductive-inference zero-shot objective described in [21], our work enhances the predictive accuracy of pretrained RS VLMs without the need of any supervision.
@@ III Method
The transductive approach employed by RS-TransCLIP is based on the hypothesis that the data structure within the feature space can be modeled as a mixture of Gaussian distributions. As a result, the RS-TransCLIP objective function integrates this hypothesis alongside affinity relationships among patches and initial text-based pseudo-labels to minimize prediction deviation. The intuition behind the proposed transductive approach is depicted in Fig. 2.
Figure 2: (a) VLMs assign each image to its closest text embedding and (b) RS-TransCLIP exploits the image-text structure to enhance the predictions without any additional labels.
@@ III-A Variable Definition
with $K$ the number of classes, $\mathcal{Q}$ the sample indices set and $\Delta_{K}$ the K-dimensional probability simplex (prediction space).
@@ III-B RS-TransCLIP objective function
GMM clustering — The goal of this term is maximizing the similarity between the assignment variables ${\mathbf{z}}_{i}$ and the likelihood ${\mathbf{p}}_{i}$. In our case, we model the likelihood of target data as a balanced mixture of $K$ multivariate Gaussian distributions. Each distribution represents a class $k$ with an associated mean vector $\boldsymbol{\mu}_{k}$ and a covariance matrix $\boldsymbol{\Sigma}$. Defining ${\mathbf{f}}_{i}\in\mathbb{R}^{d}$ as the image embedding of sample $i$, we set $p_{i,k}$ the probability that sample $i$ is generated by the Gaussian distribution of class $k$:
@@
Laplacian regularization — The aim of this term is to favor pairs of samples with high affinity to have similar assignment variables. In our case, we define non-negative affinities $w_{ij}$ using the cosine similarities between image embeddings of each sample (see line 1 in Algorithm 1). Note that affinities can be tailored for each specific use-case, provided the affinity matrix is positive semi-definite. This ensures the concavity of the term, which in turn guarantees the convergence of the decoupled updates (refer to [21] for details).
KL regularization — The purpose of this term is to prevent the assignment variables to deviate significantly from the initial pseudo-labels. In our case, we obtain the pseudo-labels $(\hat{\mathbf{y}}_{i})_{1\leq i\leq\mathcal{Q}}$ by applying the softmax function to the vector whose components are obtained by computing the dot product between the image embeddings $\mathbf{f}_{i}$ and all text embeddings $\mathbf{t}_{k}\in\mathbb{R}^{d}$, scaled by the temperature factor $\tau$ used during VLM pretraining (see line 1 in Algorithm 1). This allows us to integrate the text-knowledge into the optimization process.
@@ IV-A Experimental setup
We test RS-TransCLIP on four VLMs: CLIP [12], RemoteCLIP [11], SkyCLIP [15], and GeoRSCLIP [13] — all with various model architectures to generate their respective image embeddings. Using RS text-prompt templates from [13], 106 individual text embeddings were averaged out to get a single textual embedding per class. The zero-shot scene classification performance is evaluated on 10 RS benchmark datasets: AID, EuroSAT, MLRSNet, OPTIMAL31, PatternNet, RESISC45, RSC11, RSICB128, RSICB256, and WHURS19 [34, 35, 36, 37, 38, 39, 40, 41, 42]. Note that none of the chosen VLMs were fine-tuned on any of the listed datasets. TABLE I presents the zero-shot top-1 accuracy, without and with the addition of RS-TransCLIP.
@@ IV-B Zero-shot classification — without RS-TransCLIP
Model AID EuroSAT MLRSNet OPTIMAL31 PatternNet RESISC45 RSC11 RSICB128 RSICB256 WHURS19 Average
@@ V Conclusion
In this work, we proposed RS-TransCLIP, a transductive algorithm that enhances RS VLMs with minimal extra computational cost. By leveraging initial pseudo-labels and patch affinities, our method improves zero-shot capabilities through transductive inference, demonstrating significant accuracy improvements over inductive zero-shot classification and showing its wide applicability beyond natural images [21]. Future works will study RS-TransCLIP’s performance concerning text-prompt variability, given VLMs’ high sensitivity to input text prompts. Moreover, adapting RS-TransCLIP to a few-shot setting to incorporate labeled data will be explored in human-in-the-loop scenarios.